The cost C of printing a book is directly proportional to the number of pages p of the book. If an 36 page book costs $27 to print, how much will a 45 page book cost to print?
Answer by nerdybill(7384)

The cost C of printing a book is directly proportional to the number of pages p of the book. If an 36 page book costs $27 to print, how much will a 45 page book cost to print?
"direct proportional" think:
y = kx
solving for k:
y/x = k
.
So, given two sets of data the equation would be:
y1/x1 = y2/x2
cost1/page1 = cost2/page2
27/36 = cost2/45
27/36*45 = cost2
27/4*5 = cost2
6.75 * 5 = cost2
$33.75 = cost2